OpenStudy (anonymous):

5x + 2y = 7 y = x + 1 What is the solution set of the given system? {(5/7, 12/7)} {(12/7, 5/7)} {(9/7, 16/7)} {(16/7, 9/7)}

OpenStudy (usukidoll):

use substitution method

OpenStudy (shamim):

U hv to put the value of y frm second equation into first equation

OpenStudy (carolinar7):

Ok its better you start of with substitution

OpenStudy (carolinar7):

So solve 5x+2(x+1)=7

OpenStudy (shamim):

5x+2(x+1)=7 x=?

OpenStudy (carolinar7):

Remember Pemdas and distribute first

OpenStudy (carolinar7):

From there combine like terms and then solve for x

OpenStudy (anonymous):

i don't know what are you guys talking about

OpenStudy (shamim):

5x+2x+2=7 Right

OpenStudy (carolinar7):

7x=5

OpenStudy (shamim):

@john_97

OpenStudy (usukidoll):

divide 7 by both sides for 7x=5 then you got the x value

OpenStudy (carolinar7):

5x+2y=7 y=1+x Substitute the value for y in the first equation 5x+2(1+x)=7 Now Distribute and Combine like terms 5x+2x+2=7 -2 -2 7x=5 /7 /7 x=5/7 Now substitute the value of x in the any equation to find y

OpenStudy (arindameducationusc):

Check the attachment

OpenStudy (carolinar7):

You will get A

OpenStudy (usukidoll):

agree with @carolinar7 and @arindameducationusc